An investor has $20,000 o invest. If part is invested at 8% and the rest at 12%, how much should be invested at each rate to yield 11% on the total amount invested?
Answer by checkley77(12844)

.12X+.08(20,000-X)=.11*20,000
.12X+1,600-.08X=2,200
.04X=2,200-1,600
.04X=600
X=600/.04
X=15,000 INVESTED @ 12%
20,000-15,000=5,000 INVESTED @ 8%
PR00F:
.12*15,000+.08*5,000=2,200
1,800+400=2,200
2,200=2,200